Fermi National Laboratory


Accelerator Update for November 19, 2004 - November 22, 2004

Friday November 19
The day shift began with Linac running beam, with technicians working in the Booster, Main Injector (MI), the Antiproton Source (Pbar), and the Tevatron (TeV), and with cryogenic system techs working on the TeV sector D1 wet engine.

Operations reported at 4:58 PM that work in Booster included: replacing a power supply, repairing a BPM, and fixing a collimator motion controller.

Around 5 PM, operators began interlocking the Collider Detector at Fermilab (CDF) collision hall and the MI tunnel.

Operators turned Booster on at 6 PM and around 8 PM ran a short amount of beam and then shut it off.

Operators completed interlocking D-Zero (D0) at 8:08 PM.

During the TeV turn on, at 9:05 PM, operators discovered a VCB stuck open at TeV sector E3. Experts came in and got the breaker to close, but think there is still a wire connection problem.

Saturday November 20
The midnight shift began with Booster on, but with no beam, with MI in access, and with the TeV turn on underway.

D0 experimenters reported at 12:21 AM that they turned on their solenoid.

At 1:15 AM, operators reported trouble with the main bus QBS ramp. With expert help, QBS began ramping normally.

Operations had TeV at its first 980 GeV ramp at 1:54 AM.

At 2:02 AM a TeV power supply, B0Q3, tripped off.

Operations turned off the TeV for an access at 4:21 AM.

An MI expert reported at 11:43 AM that their bake was progressing normally and should be completed around 4 PM.

A Booster expert began running beam in Booster at 12:53 PM. About an hour and a half later, he handed off tuning to Operations.

The alignment crew was finished with their work on the TeV F17 kicker and out of the tunnel by 3:24 PM. Vacuum work continued until 6 PM.

Cryo techs put a "band aide" on the TeV A4 helium leak and then started an A4 cooldown at 3:30 PM.

Operators interlocked MI at 7:15 PM and had a ramp established by 8:10 PM.

Operations reported the TeV resumed ramping at 9:40 PM.

At 11:01 PM, an A4 lead failure caused the ramp to dump.

Sunday November 21
The midnight shift began with Linac and Booster running beam, and with operators testing Switchyard power supplies.

Operators interlocked D0 at 12:23 AM.

Operators dumped the ramp at 1:07 AM because the A4 lead flow was about to trip anyway. Operators adjusted the lead flow and started turning the TeV back on at 1:42 AM.

Operators made D0 ready for an access by 12:15 PM.

Vacuum techs reported at 2 PM that they discovered a problem with the bake at MI-30. Operations informed an expert.

Vacuum techs came out of MI and operators resumed ramping at 5:11 PM.

At 7 PM, Operations reported the following events: Experts began MI power supply measurements, an expert began working with the P1 line power supplies, an expert started conditioning the MI-52 slow extraction septa, the TeV was in a dry squeeze, and D0 experimenters were accessing their collision hall.

Monday November 22
The midnight shift began with an expert conducting power supply studies in the P1, P2, A1, and A2 transfer lines.

Operators began preparing MI for access at 4:34 AM.

Operators had CDF prepared for access by 6:16 AM.

Operators had the TeV prepared for access by 8 AM.

Plans for Monday
The plan is to finish necessary work in all machines so that we can start running beam in the TeV by Monday, November 29.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/23/2004   email Fermilab